1. What's the future value of $55,000 after 20 years if the appropriate interest rate is 3%, compounded semiannually?

2. Tucson Bank offers to lend you $50,000 at a nominal rate of 12%, compounded monthly. The loan (principal plus interest) must be repaid at the end of the year. Phoenix Bank also offers to lend you the $50,000, but it will charge an annual rate of 10.8%, with no interest due until the end of the year. How much higher or lower is the effective annual rate charged by Tucson versus the rate charged by Phoenix?

3. You are offered a chance to buy an asset for $200,500 that is expected to produce cash flows of $100,000 at the end of Year 1, $42,000 at the end of Year 2, $52,850 at the end of Year 3, and $43,250 at the end of Year 4. What rate of return (IRR) would you earn if you bought this asset?
Answer #1

Future value = Present value X (1+i=n)^nt Future value after 20 years =fv = $99,771.01 $55,000 X (1+3%/2)^(2*20)